OpenStudy (anonymous):

Solve the equation by completing the square. x² - 2x = 15

hero (hero):

Hint: add (-2/2)^2 to both sides

OpenStudy (anonymous):

x^2-2x-15=0 x^2-5x+3x-15=0 x(x-5)+3(x-5)=0 (x+3)(x-5)=0 x=-3 or x=5